Community Coffee Co. wants a new flavor of Cajun coffee. How many pounds of coffee worth $7 a pound should be added to 14 pounds of coffee worth $4 a pound to get a mixture worth $5 a pound?
Answer by jorel1380(3719)

7x+14(4)=5(x+14)
7x+56=5x+70
2x=14
x=7
You need 7lbs. of $7.00 coffee for the required mix..